Determine the partial pressure (in mmHg) of each gas in a container with:

The total pressure is 850 mmHg

2.0 moles of N2

3.0 moles of O2

7.0 moles of H2
1
Expert's answer
2020-04-18T07:00:25-0400

Since, by the condition of the question, we don't know the value of temperature, but since all gases are contained in one vessel, it's the same for all. So we can write the equivalences:

"T = (p(1) • V)/(\nu(1) • R) = (p(2) • V)/(\nu(2)•R) = (p(3)•V)/(\nu(3)•R) = (p•V)/(\nu•R)", also V (for this system) and R are constant too, so we can express "V\/R" as constant k:

"(p(1)•k)/\nu(1) = (p(2)•k)/\nu(2) = (p(3)•k)/\nu(3) = (p•k)/\nu" and now we can divide all parts of the equivalence by the constant k:

"p(1)/\nu(1) = p(2)/\nu(2) = p(3)/\nu(3) = p/\nu" = 850(mmHg)/(2+3+7)(mol) = 70.833(mmHg/mol)

So the partial pressure will be calculated: "p(1) = 70.833(mmHg\/mol)\u20222(mol) =" 141.67 mmHg - for N2

"p(2) = 70.833(mmHg)\u20223(mol) =" 212.50 mmHg - for O2

"p(3) = 70.833(mmHg )\u20227(mol)" = 495.83 mmHg - for H2

Answer: p(N2) = 141.67 mmHg; p(O2) = 212.50 mmHg; p(H2) = 495.83 mmHg.